OpenStudy (anonymous):

Solve the equation -x/3 = 27 & 5c-9=8-2c Help?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

Is that all you need to do for the first one?

OpenStudy (anonymous):

Oooooooh

OpenStudy (anonymous):

Thank you thats a lot easier than i thought

OpenStudy (anonymous):

How do i solve? I'm slightly confused do i need to subtract?

OpenStudy (texaschic101):

do you understand any of these problems....I mean about how to do them ?

OpenStudy (anonymous):

Yeah second one! I really don't understand how to go about solving equations like this.

OpenStudy (texaschic101):

even the first one ? Do you understand how it was done ?

OpenStudy (anonymous):

The first makes sense

OpenStudy (texaschic101):

5c - 9 = 8 - 2c ... we want to get the c terms on one side and everything else on the other side...and what you do on one side of the equal sign, you have to do on the other side. So first we will add 2c to both sides 5c - 9 + 2c = 8 - 2c + 2c -- now we combine like terms 7c - 9 = 8 ...see how the 2c's on the right side cancelled out now we will add 9 to both sides 7c - 9 +9 = 8 + 9 -- combine like terms 7c = 17....now we need c by itself, so we divide both sides by 7 (7/7)c = 17/7....the 7/7 turns into 1 and 1c is just c.... so c = 17/7 or in decimal form = 2.43

OpenStudy (anonymous):

Thank you!! im taking notes

OpenStudy (anonymous):

did u solve the second one?